# A deep backward regression-based scheme for high-dimensional nonlinear partial differential equations

### 16 Mar 2026 v1

@@ 3.2 Error analysis
We now estimate the error of the numerical solutions $(\mathcal{Y}_{i},\mathcal{Z}_{i})_{0\leq i\leq N-1}$ and the analytical solutions $(Y_{t_{i}},Z_{t_{i}})_{0\leq i\leq N-1}$. This result provides an error estimate of the DBR method (3.9) in terms of the $L^{2}$-sense.
The standard assumptions which guarantee the existence and uniqueness on the coefficients of the FBSDEs (2.1) associated to the semilinear PDEs (1.1) are made as following:
@@
for all $t\in[0,T],$ $x_{1},x_{2}\in\mathbb{R}^{d}$, $y_{1},y_{2}\in\in\mathbb{R}^{n}$ and $z_{1},z_{2}\in\mathbb{R}^{n\times d}$; and suppose
@@
Now, we investigate the errors of the scheme (3.9) and define, for $i=N-1,N-2,\cdots,0$
$\left\{\begin{aligned} \mathcal{U}_{t_{i}}^{m}&:=\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})+hf\left(t_{i},X_{i}^{\pi,m},\mathcal{U}_{t_{i}}^{m},\mathcal{V}_{t_{i}}^{m}\right)\right],\\ \mathcal{V}_{t_{i}}^{m}&:=\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\frac{(\Delta W_{i}^{m})^{\top}}{h}\right].\end{aligned}\right$. (3.10)
Theorem 3.1
Suppose the assumptions (i)-(iv) hold. Let $(Y_{t_{i}},Z_{t_{i}})$ and $(\mathcal{Y}_{i},\mathcal{Z}_{i})$ be solutions of the FBSDEs (2.1) and solutions of the DBR method (3.9) respectively. Then, we have, for small enough $h$
@@
$\quad+C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\left|R_{i+1,2}^{y}\right|^{2}+\left|R_{i,2}^{z}\right|^{2}+N\left|R_{i+1,1}^{y}\right|^{2}+\left|R_{i,1}^{z}\right|^{2}\right],$
@@
where $C$ represents a positive generic constant which is independent of $\pi$ and may change from line to line.
@@
Proof. Step 1. On the uniform discrete mesh $\pi$, we can write the BSDE in (2.1) at the mesh points $t_{i}$ as follows
$Y_{t_{i}}=Y_{t_{i+1}}+\int_{t_{i}}^{t_{i+1}}f_{s}ds-\int_{t_{i}}^{t_{i+1}}Z_{s}dW_{s},$ (3.11)
where $f_{s}=f(s,X_{s},Y_{s},Z_{s})$. From (3.10) and (3.11), we have, for $i\in\{0,1,\ldots,N-1\}$
$Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}=$ $\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]$ (3.12)
$+\mathbb{E}_{i}\left[\int_{t_{i}}^{t_{i+1}}\left(f_{s}-f\left(t_{i},X_{i}^{\pi,m},\mathcal{U}^{m}_{t_{i}},\mathcal{V}_{t_{i}}^{m}\right)\right)ds\right].$
From Young inequality in the form: $(a+b)^{2}\leq(1+\gamma h)a^{2}+(1+\frac{1}{\gamma h})b^{2},$ for $a,b\in\mathbb{R}$, and some $\gamma>0$ which will be given later, Cauchy-Schwarz inequality, and the Lipschitz condition with respect to $f$, we obtain
$\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]$
$\leq(1+\gamma h)\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\quad+(1+\frac{1}{\gamma h})\mathbb{E}\left[\left|\mathbb{E}_{i}\left[\int_{t_{i}}^{t_{i+1}}\left(f_{s}-f\left(t_{i},X_{i}^{\pi,m},\mathcal{U}^{m}_{t_{i}},\mathcal{V}^{m}_{t_{i}}\right)\right)ds\right]\right|^{2}\right]$
$\leq(1+\gamma h)\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\quad+h(1+\frac{1}{\gamma h})\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}\left(f_{s}-f\left(t_{i},X_{i}^{\pi,m},\mathcal{U}^{m}_{t_{i}},\mathcal{V}^{m}_{t_{i}}\right)\right)^{2}ds\right]$
$\leq(1+\gamma h)\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\quad+4L^{2}h(1+\frac{1}{\gamma h})\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}\left(|s-t_{i}|+|X_{s}-X_{i}^{\pi,m}|^{2}+|Y_{s}-\mathcal{U}^{m}_{t_{i}}|^{2}+|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}\right)ds\right].$ (3.13)
Because the functions $\mu$ and $\sigma$ are the Lipschitz continuity (see the assumptions (i)), we have (see [30])
$\max_{0\leq i\leq N-1}\mathbb{E}\left[\sup_{s\in\left[t_{i},t_{i+1}\right]}\left|X_{s}-X_{i}^{\pi,m}\right|^{2}\right]\leq Ch,$ (3.14)
$\mathbb{E}\left[\sup_{0\leq t\leq T}\left|X_{t}\right|^{p}\right]\leq C_{p}(1+|x_{0}|^{p}),\quad p\geq 1,$ (3.15)
where $C_{p}$ is a constant depending on $p,\mu,\sigma$ and $T$. From the assumptions (ii)-(iii) and (3.15), we have (see [19, 36])
$\max_{0\leq i\leq N-1}\mathbb{E}\left[\sup_{t\in\left[t_{i},t_{i+1}\right]}\left|Y_{t}-Y_{t_{i}}\right|^{2}\right]$ $\leq Ch,$ (3.16)
$\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}ds\right]$ $=\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+h\mathbb{E}\left[|\bar{Z}_{t_{i}}-\mathcal{V}^{m}_{t_{i}}|^{2}\right],$ (3.17)
$\mathbb{E}\left[\int_{0}^{T}f_{s}^{2}ds\right]$ $\leq+\infty,$ (3.18)
where $\bar{Z}_{t_{i}}=\mathbb{E}_{i}\left[\frac{1}{h}\int_{t_{i}}^{t_{i+1}}Z_{t}\mathrm{~d}t\right]$. From (3.14), (3.16) and (3.17), we can rewrite (3.13) as
$\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]\leq$ $(1+\gamma h)\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$ (3.19)
$+4L^{2}h(1+\frac{1}{\gamma h})\left(Ch^{2}+2h\mathbb{E}\left[|Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}|^{2}\right]\right)$
$+4L^{2}h(1+\frac{1}{\gamma h})\left(\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+h\mathbb{E}\left[|\bar{Z}_{t_{i}}-\mathcal{V}^{m}_{t_{i}}|^{2}\right]\right).$
Multiplying (3.11) with $\Delta W_{i}^{\top}$ and then taking the conditional expectation on both sides of the derived equation, we obtain
$0=\mathbb{E}_{i}[Y_{t_{i+1}}\Delta W_{i}^{\top}]+\mathbb{E}_{i}\left[\Delta W_{i}^{\top}\int_{t_{i}}^{t_{i+1}}f_{s}ds\right]-\int_{t_{i}}^{t_{i+1}}\mathbb{E}_{i}[Z_{s}]ds.$ (3.20)
From (3.10) and (3.20), we have, for $i\in\{0,1,\ldots,N-1\}$
$\bar{Z}_{t_{i}}-\mathcal{V}^{m}_{t_{i}}=$ $\mathbb{E}_{i}\left[\left(Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right)\frac{\Delta W_{i}^{\top}}{h}\right]+\mathbb{E}_{i}\left[\frac{\Delta W_{i}^{\top}}{h}\int_{t_{i}}^{t_{i+1}}f_{s}ds\right]$
$=$ $\mathbb{E}_{i}\left[\left(Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})-\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right)\frac{\Delta W_{i}^{\top}}{h}\right]$
$+\mathbb{E}_{i}\left[\frac{\Delta W_{i}^{\top}}{h}\int_{t_{i}}^{t_{i+1}}f_{s}ds\right].$
$\mathbb{E}\left[\left|\bar{Z}_{t_{i}}-\mathcal{V}^{m}_{t_{i}}\right|^{2}\right]\leq$ $\frac{2d}{h}\mathbb{E}\left[\left|Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right|^{2}\right]$ (3.21)
$-\frac{2d}{h}\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right|^{2}\right]+2d\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}f_{s}^{2}ds\right].$
Plugging (3.21) into (3.19), we have
$\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]\leq$ $(1+Ch)\mathbb{E}\left[\left|Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right|^{2}\right]+Ch\mathbb{E}\left[|Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}|^{2}\right]+Ch^{2}$
and hence for small enough $h$, we have
$\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]\leq$ $(1+Ch)\mathbb{E}\left[\left|Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right|^{2}\right]+Ch^{2}$ (3.22)
@@
$\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]\geq$ $(1-h)\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]-\frac{1}{h}\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right].$ (3.23)
Inserting (3.23) into (3.22), we have
$\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]$ $\leq(1+Ch)\mathbb{E}\left[\left|Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right|^{2}\right]+Ch^{2}$ (3.24)
$\quad+C\left(\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+h\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}f_{s}^{2}ds\right]+\frac{1}{h}\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]\right).$
With the help of the discrete Gronwall’s lemma and (3.18), we derive
$\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]$ $\leq C\mathbb{E}\left[\left|\Phi(X_{T})-\Phi(X_{N}^{\pi,m})\right|^{2}\right]+Ch$ (3.25)
$\quad+C\left(\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+\frac{1}{h}\sum\limits_{i=0}^{N-1}\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]\right).$
@@
$\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]\leq$ $2\mathbb{E}\left[\left|\mathcal{Y}_{i}-S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)\right|^{2}\right]+2\mathbb{E}\left[\left|S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right].$ (3.26)
In what follows, we handle each term in (3.26) separately.

### 21 May 2026 v3

@@ 3.2 Error analysis
We now estimate the error of the numerical solutions $(\mathcal{Y}_{i},\mathcal{Z}_{i})_{0\leq i\leq N-1}$ and the analytical solutions $(Y_{t_{i}},Z_{t_{i}})_{0\leq i\leq N-1}$. This result provides an error estimate of the DBR method (19) in the $L^{2}$ sense.
The standard assumptions which guarantee the existence and uniqueness on the coefficients of the FBSDEs (5) associated to the semilinear PDEs (1) are made as following:
@@
for all $t\in[0,T],$ $x_{1},x_{2}\in\mathbb{R}^{d}$, $y_{1},y_{2}\in\mathbb{R}^{n}$ and $z_{1},z_{2}\in\mathbb{R}^{n\times d}$; and suppose
@@
In the error analysis below, the same time-step condition (10) is imposed. Consequently, all implicit equations appearing in the DBR scheme and in the auxiliary processes are well-defined.
Now, we investigate the errors of the scheme (19) and define, for $i=N-1,N-2,\cdots,0$
$\left\{\begin{aligned} \mathcal{U}_{i}^{m}&:=\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})+hf\left(t_{i},X_{i}^{\pi,m},\mathcal{U}_{i}^{m},\mathcal{V}_{i}^{m}\right)\right],\\ \mathcal{V}_{i}^{m}&:=\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\frac{(\Delta W_{i}^{m})^{\top}}{h}\right],\end{aligned}\right$. (21)
and by the Markov property of the discretized forward process $\{X_{i}^{\pi,m}\}_{0\leq i\leq N}$, we have
$\mathcal{U}_{i}^{m}=\widetilde{u}(t_{i},X^{\pi,m}_{i}),\quad\mathcal{V}_{i}^{m}=\widetilde{v}(t_{i},X^{\pi,m}_{i}),$ (22)
where $\widetilde{u}$ and $\widetilde{v}$ are some deterministic (but unknown) functions;
$\left\{\begin{aligned} \widetilde{\mathcal{U}}_{i}^{m}&:=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m,k},\theta^{*}_{y,i+1})+hf\left(t_{i},X_{i}^{\pi,m},\widetilde{\mathcal{U}}_{i}^{m},\widetilde{\mathcal{V}}_{i}^{m}\right),\\ \widetilde{\mathcal{V}}_{i}^{m}&:=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m,k},\theta^{*}_{y,i+1})\frac{(\Delta W_{i}^{m,k})^{\top}}{h}.\end{aligned}\right$. (23)
Let us introduce
$\varepsilon_{i}^{y}=\inf\limits_{\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}}\mathbb{E}\left[\left|\widetilde{u}(t_{i},X^{\pi,m}_{i})-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta^{r_{j}}_{y,i})\right|^{2}\right],$ (24)
$\varepsilon_{i}^{z}=\inf\limits_{\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}}\mathbb{E}\left[\left|\widetilde{v}(t_{i},X^{\pi,m}_{i})-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta^{r_{j}}_{z,i})\right|^{2}\right],$
and
$\widetilde{\varepsilon}_{i}^{y}=\inf\limits_{\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}}\mathbb{E}\left[\left|\widetilde{\mathcal{U}}_{i}^{m}-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta^{r_{j}}_{y,i})\right|^{2}\right],$ (25)
$\widetilde{\varepsilon}_{i}^{z}=\inf\limits_{\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}}\mathbb{E}\left[\left|\widetilde{\mathcal{V}}_{i}^{m}-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta^{r_{j}}_{z,i})\right|^{2}\right],$
By a slight abuse of notation, we henceforth write $\widetilde{\mathcal{V}}_{i}^{m}$ to denote $S_{z,i}\left(p_{i}^{m},\Delta W_{i}^{m},X_{i+1}^{\pi,m}\right)$, to shift the focus from the specific functional form of the network to its role as an approximated function.
Population-loss convention. The loss functions in (17) are empirical losses used for the implementation of Algorithm 1. In the following error analysis, however, we work with their population counterparts. More precisely, the expectation $\mathbb{E}[\cdot]$ is taken with respect to all sources of randomness involved in $X_{i}^{\pi,m}$, $\widetilde{\mathcal{U}}_{i}^{m}$, and $\widetilde{\mathcal{V}}_{i}^{m}$. For fixed $i$ and $r_{j}$, we define
$\mathcal{F}_{z,i}^{r_{j}}(\theta_{z,i}^{r_{j}})$= $\mathbb{E}\left[\left|\widetilde{\mathcal{V}}_{i}^{m}-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{r_{j}})\right|^{2}\right],$ (26)
and, after fixing the population minimizer $\theta_{z,i}^{*}$,
$\mathcal{F}_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})$= $\mathbb{E}\Bigg[\bigg|\widetilde{\mathcal{U}}_{i}^{m}+h\Big(f\big(t_{i},X_{i}^{\pi,m},\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}}),\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*})\big)$ (27)
$\qquad\qquad-f\big(t_{i},X_{i}^{\pi,m},\widetilde{\mathcal{U}}_{i}^{m},\widetilde{\mathcal{V}}_{i}^{m}\big)\Big)-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\bigg|^{2}\Bigg].$
Throughout the proof of Theorem 3.1, the symbols $F_{y,i}^{r_{j}}$ and $F_{z,i}^{r_{j}}$ are understood as the population losses $\mathcal{F}_{y,i}^{r_{j}}$ and $\mathcal{F}_{z,i}^{r_{j}}$, respectively. The empirical losses in (17) are Monte Carlo approximations of these population losses. Therefore, Theorem 3.1 should be interpreted as an approximation-error estimate under exact minimization of the population losses. The additional errors arising from replacing the population losses by their empirical counterparts and from non-exact stochastic optimization are not included in this theorem; they can be incorporated separately as generalization and optimization residuals.
Theorem 1.
Suppose the assumptions (i)-(iv) hold. Let $(Y_{t_{i}},Z_{t_{i}})$ and $(\mathcal{Y}_{i},\mathcal{Z}_{i})$ be solutions of the FBSDEs (5) and solutions of the DBR method (19) respectively. Then, for $h$ small enough and satisfying (10), we have
@@
$\quad+C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\varepsilon_{i}^{y}+\varepsilon_{i}^{z}+\frac{N}{K}\right],$
@@
where $K=\min\limits_{0\leq i\leq N-1}\{K_{i}\}$; $\bar{Z}_{t_{i}}=\mathbb{E}_{i}\left[\frac{1}{h}\int_{t_{i}}^{t_{i+1}}Z_{t}\mathrm{~d}t\right]$; $C$ represents a positive generic constant independent of $\pi$, which may depend on $d$ and change from line to line.
Proof.
@@
Step 1. On the uniform discrete mesh $\pi$, we can write the BSDE in (5) at the mesh points $t_{i}$ as follows
$Y_{t_{i}}=Y_{t_{i+1}}+\int_{t_{i}}^{t_{i+1}}f_{s}ds-\int_{t_{i}}^{t_{i+1}}Z_{s}dW_{s},$ (28)
where $f_{s}=f(s,X_{s},Y_{s},Z_{s})$. From (21) and (28), we have, for $i\in\{0,1,\ldots,N-1\}$
$Y_{t_{i}}-\mathcal{U}^{m}_{i}=$ $\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]$ (29)
$+\mathbb{E}_{i}\left[\int_{t_{i}}^{t_{i+1}}\left(f_{s}-f\left(t_{i},X_{i}^{\pi,m},\mathcal{U}^{m}_{i},\mathcal{V}_{i}^{m}\right)\right)ds\right].$
Following a similar proof strategy for the convergence of the Euler method in [2, 37], we have
$\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{U}^{m}_{i}\right|^{2}\right]\leq$ $(1+Ch)\mathbb{E}\left[\left|Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right|^{2}\right]+Ch^{2}$ (30)
@@
$\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{U}^{m}_{i}\right|^{2}\right]\geq$ $(1-h)\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]-\frac{1}{h}\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{i}\right|^{2}\right].$ (31)
Inserting (31) into (30), we have
$\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]$ $\leq(1+Ch)\mathbb{E}\left[\left|Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right|^{2}\right]+Ch^{2}$ (32)
$\quad+C\left(\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+h\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}f_{s}^{2}ds\right]+\frac{1}{h}\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{i}\right|^{2}\right]\right).$
With the help of the discrete Gronwall’s lemma, we derive
$\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]$ $\leq C\mathbb{E}\left[\left|g(X_{T})-g(X_{N}^{\pi,m})\right|^{2}\right]+Ch$ (33)
$\quad+C\left(\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+\frac{1}{h}\sum\limits_{i=0}^{N-1}\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{i}\right|^{2}\right]\right).$
@@
$\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{i}\right|^{2}\right]\leq$ $2\mathbb{E}\left[\left|\mathcal{Y}_{i}-\widetilde{\mathcal{U}}_{i}^{m}\right|^{2}\right]+2\mathbb{E}\left[\left|\widetilde{\mathcal{U}}_{i}^{m}-\mathcal{U}^{m}_{i}\right|^{2}\right].$ (34)
In what follows, we handle each term in (34) separately.